Cambonesia is a small exporter of bananas. Without trade, the price for bananas is $1,200 per tonne. The world price of bananas is $2,000 per ton. Currently, Cambonesian exporters pay $500 for every tonne of bananas shipped abroad. Suppose that the government of Cambonesia decides to pay domestic producers half of the transportation costs for every tonne of banana exported. Discuss the welfare implications of such a policy in the market for bananas in Cambonesia. Illustrate your answer with appropriate diagrams.

(Hint: You should consider the impact on consumer surplus, producer surplus, the government’s revenue or expenditure, and deadweight loss (if any), within the market.)
Answer #1

In the diagram given below, we have shown the demand and supply of bananas for a small country, Cambonesia .

Price Supply $2000 $1700 $1450 $1200 World price Domestic price + export duty -Domestic price + half of export duty Domestic

Domestic price of bananas in Cambonesia is $1,200 per ton but the world price of bananas is $2,000 per ton. Thus, it will trade and export bananas.

A shipping charge or export duty of $500 per ton is applied. This is the source of revenue for government. Thus, the export price becomes $1,700 per ton. Now, at this price, quantity Q5 of bananas can be supplied but a quantity of only Q2 is demanded.

Thus, at export price of $1,700, consumer surplus = Area (a+b).

Government revenue = Area (c+d)

Producer surplus = Area e

Dead weight loss = Area (g+h)

Thus, total welfare = Consumer surplus + government revenue + producer surplus - dead weight loss = Area (a+b+c+d)-Area (g+h) ....(i)

Now, if shipping charge is halved to $250, export price becomes $1,450. At this price, quantity demanded is Q3 whereas quantity that could be supplied is Q4.

Thus, at export price of $1,450 , consumer surplus = Area (a+b+c)

Government revenue = Area d

Producer surplus = Area e

Dead weight loss = Area h

Thus, total welfare = Consumer surplus + government revenue + producer surplus - dead weight loss Area (a+b+c+d+e) - Area h...(ii)

Comparing (i) and (ii), we can say , that decreasing export prices have negative impact on total welfare.